The Capital by Karl Marx

by Hadas Zeilberger

Trade is a very important procedure that society cannot do without but before one gets there they need distinct values for commodities. In Chapter One of Capital by Karl Marx, Marx starts off with discussing the values of commodities. The main values are Use-value, Quantitative- value, and Qualitative-value. Getting more specific, there’s exchange-value, relative-value, and labor value.

The usefulness of a commodity is referred to as its use-value. Since this is the measurement of how useful it is, and the usefulness of an object is its sole purpose, use-value makes us the physical body of the commodity. Everything with use-value can be looked at from two points of view; qualitatively and quantitatively. The quality of an object refers to the many properties that make it up. It can, therefore, be used in a number of ways, some of which have not yet been discovered. Quantitative value is usually used with the exchange value of a commodity.

Exchange value is when use-value(s) is exchanged with other use-value(s). The value of the object is purely relative to the other objects it is being exchanged with, because it depends on the circumstances around it. For example, if a certain number of x is exchangeable with a different number of y, z, and v, then x’s value differs with each thing it is exchanged with, but y, z, and v must all have the same magnitude of value, the value of a certain number of x. Another circumstance of exchange is if there are two commodities, A and B. A certain number of A has to be equal to a number of B. Once A = B, A and B must be equal to a third commodity, which we will call C. Labor-Value is also considered in giving a commodity an exchange value.

Labor-Value is the labor that goes behind making a commodity. It can completely change an exchange value of a useful thing, depending on how hard the labor is. For example, if it takes twice as long to produce an ear of corn than usual, than one ear of corn usually equals one gallon of milk, then the new exchange value for the milk is one gallon milk = 1/2 ear of corn. If the labor that goes into producing milk doubles, then the corn raises in exchange value to one ear of corn = 1/2 gallon of milk. Now let’s say that milk and corn both vary slightly in labor-value. Then the exchange value stays the same, which is one ear of corn = one gallon of milk. This form is called the relative form. When using labor to make use-value, it is referred to as Use-Labor-Value. The relative form is a very concrete form of labor-value because it gives an equation of the exact amount of time to produce one object with the exact time of another and compares/varies them. When going into a more abstract form, however, the same kind of labors simply equal each other if they take the same amount of time. But, one labor will produce a larger quantity of the commodity than the other. Therefore, the labor that produced a bigger quantity of use-values is now cut down on the value if the total number of objects.

Marx goes into much detail concerning the values of commodities, talking about different forms beside use-values. Qualitative and quantitative forms, exchange value, and labor value are the most important and major values. Marx’s system of trading is ingenious and has revolutionized the world’s economics.